OpenStudy (anonymous):

Can someone explain this? Find the derivative of f(x) = 12x^2 + 8x at x = 9.

OpenStudy (welshfella):

the formula for the derivative of an algebraic function like f(x) = a x^n is f'(x) = an x^(n-1)

OpenStudy (welshfella):

so for example f(x) = 4x^3 f'(x) = 12 x^(3-1) = 12 x^2

OpenStudy (welshfella):

so find the derivative of 12x^2 and 8x then replace x by 9 to find f'(2)

OpenStudy (welshfella):

* f'(9)

OpenStudy (welshfella):

the derivative gives the slope of the tangent to the curve at x = 9

OpenStudy (anonymous):

so I have f(x) = 12x^2 + 8x at x = 9. and you say that a = 12 n = 2 so 24x^1?

OpenStudy (anonymous):

and what about the +8x

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

You must get the derivative of 8x aswell and add it to the other part of the derivative :)